the bond price inorease or decrease from its current market value? Explain why (No caloulations required b wir the price be piice be higher or lower than $1,000? Explain why. (No calculations required) you hokd be your total return in dollars and percentage? the bond for the entire year and the change in yield to maturity ocoure, what would Problem& 20 points) tast is to vailue the stock price of Mare Co. with the Dvidend Growth Model (DGM) in constant growth. You have the following information: $4.50 Recent dividend per share (Ds) Risk-free rate (Rrf) Beta of the stock (B) Average stock return on the market (Rm) Estimated long-term dividends growth rate (g) 3.25% 1.60 11% 5% 1) Calculate the value of the stock of Mare Co using the Dividend Growth Model (DGM) in constant growth. 2) The stock currently trades at $34.86 in the stock market: would you recommend buying it? www.euruntedu
Answer #1
3..
Value of the stock is the sum of the present values of the future dividends on the stock.
So, first, we need to find the required rate of return on the stock, with which to discount the future dividend-cashflows on the stock , to their present values--
using CAPM formula,
Required Return on Equity=Risk-free Rate+(Beta*Market risk Premium)
where,
Market Risk premium=Market return on the stock-Risk-Free Rate
Plugging in the given values in the above formula,
we get the
Required Return on Equity=3.25%+(1.6*(11%-3.25%))=
15.65%
Now, as per Dividend Growth Model(DGM)
Cost of Equity or Reqd. return on Equity=(Next Dividend/Current stock price)+Growth rate of dividends
where,
Next dividend=Current dividend*(1+Growth rate of dividends)
ie. 4.5*(1+5%)=
4.725
Plugging in the rest of the values,but for stock price,
15.65%=(4.725/Current stock Price)+5%
Solving the above,
the value of the stock,ie. The current stock price=44.37
Alternately,re-stating the above formula,
Value of the stock=Next dividend/(Reqd. return-Growth rate of dividends)
ie. Value of the stock=D1/(r-g)
ie.4.725/(15.65%-5%)=
44.37
2..
Value of the stock as per DGM            = $ 44.37
Currently trading in the market at    =   $ 34.86
The stock is under-valued in the market .
Indicates low demand---or less investor demand.
Hence not recommended for buying/investing.